# PeersimGym: An Environment for Solving the Task Offloading Problem with Reinforcement Learning

### v1

@@ 1 Introduction
The proliferation of large networks of devices, such as the Internet of Things (IoT), has led to an exponential increase in data generation, requiring significant computational resources for processing. Traditionally, Cloud Computing provided the backbone for such computational demands. However, its limitations in latency and network traffic have become apparent with the growth of device networks [1]. Edge Computing emerged as a paradigm shift, extending the Cloud to bring processing capabilities closer to data sources, mitigating latency and traffic issues. In this evolving landscape, task offloading—the distribution of computational tasks across network participants—has gained prominence, particularly within the Fog and Multi-access Edge Computing (MEC) paradigms. Both paradigms aim to decentralize computing power, bringing it closer to end-users and alleviating the constraints of traditional Cloud Computing [2, 3]. Despite distinctions between MEC and Fog computing, this paper treats them interchangeably, focusing on their shared goal of minimizing device-to-cloud distances [4]. Addressing the challenges of task offloading in such distributed environments involves balancing numerous factors, including task latency, energy consumption, and task completion reliability [5]. Traditional optimization techniques often fall short in efficiently managing these complex systems, paving the way for Reinforcement Learning (RL), specifically, Multi-Agent Reinforcement Learning (MARL) as promising solutions. The ability of MARL agents to iteratively learn optimal strategies through environment interaction makes them particularly suited for decentralized edge systems [6, 7].
Given the impracticality of training agents in live systems, simulation emerges as a vital tool. Yet, the lack of a standardized, customizable simulation platform has been a bottleneck, compelling researchers to develop bespoke, often suboptimal, solutions for their experimental setups.
@@
Contributions. Our contributions are focused on advancing the field of MARL and Edge computing through the introduction of the PeersimGym environment and comprehensive experimental analysis. The PeersimGym Environment is a highly adaptable simulation platform tailored for the development, training, and evaluation of MARL strategies for task offloading challenges in Edge Computing systems. This environment, built on the synergy between the Peersim simulator and the PettingZoo API [8], allows detailed configuration of network topologies, node characteristics, and task parameters, facilitating a wide range of experimental setups. An experimental analysis demonstrates the capability of the PeersimGym to train effective MARL solutions, exemplified by the performance of a Double Deep Q Network and Advantage Actor Critic algorithm. Our analysis compares the performance of DRL agents to several non-RL algorithms across various network configurations and task offloading scenarios, highlighting the advantages of RL approaches in optimizing Edge Computing networks.
@@
The source code for PeersimGym, along with usage documentation and testing resources, is made available in the Simulator repository, and the Agent repository, fostering further research and community engagement in advancing Edge Computing solutions.
@@ 3 Enhancing Task Offloading with MARL
To address the complexity of Task Offloading in highly complex edge environments, researchers leverage RL and its subset, Deep Reinforcement Learning (DRL), and by exploring the distributed nature of these systems MARL emerges as promising solution. As a tool to develop MARL algorithms, PeersimGym incorporates a Python-based environment tailored for developing, training, and deploying RL models for task offloading, aligned with the PettingZoo framework. With the set of experiments in this paper, we showcase how our simulator can enable further RL and DRL contributions to this field 5.
Certain Nodes within the network act as RL agents. These agents have the capability to observe the state (albeit partially) and make informed decisions regarding whether to process tasks locally or offload them. This decision-making process is influenced by the need to balance between local processing, the risk of queue overflow, and the costs associated with task offloading, including potential overload of other nodes.
@@
Deep Reinforcement Learning for Task Offloading. Deep Q Networks (DQN) [12] and their variants, such as Double Deep Q Networks (DDQN) [25] and Actor-Critic methods, like the Advantage Actor-Critic (A2C)[11], are prominent DRL approaches applied to task offloading. These methodologies have been shown to stabilize training and enhance learning efficacy through sophisticated neural network architectures that approximate optimal action-selection policies.
In PeersimGym, the flexibility of our RL environment supports the integration of various DRL models. For illustrative purposes, we focus on implementations utilizing DDQN and A2C, reflecting their proven effectiveness in recent literature [13, 11]. This choice underscores the potential of DRL to address the complexities associated with task offloading in Edge Computing, as shown in 5.
Framework and Model Dynamics. Task offloading decisions within our simulated environment are modeled as a Markov Game (MG), accommodating the multi-agent aspect of Edge Computing networks. This formulation extends the Markov Decision Process (MDP) framework to scenarios involving multiple decision-makers, thereby capturing the interactive and competitive nature of task offloading. Single-agent RL is modeled as a Markov Decision Process (MDP), with a sequence of states such that the Markov Property holds, i.e., the next state $s_{t+1}$ depends exclusively on the current state $s_{t}$ and the performed action $a_{t}$. When including multiple agents, most of the MDP convergence properties do not hold, therefore, we must formulate our problem as a Markov Game (MG) [26]. In our setting, in a network of $N$ nodes, an MG is represented as a tuple $\langle n,\mathcal{S},\mathcal{A},P,R,\gamma\rangle$, where $n$ is the number of agents (i.e., nodes with a controller protocol), $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $P$ is an (unknown) transition probability function, $R$ is the reward function, and $\gamma$ is a reward discount factor. Below, we introduce these constructs in more detail.
State Space. The state space in our environment is designed to be highly customizable, enabling the modeling of various Edge Computing scenarios. It encompasses node IDs, their tiers(i.e. edge, fog, cloud), queue sizes, processing capabilities, positions, bandwidths, and transmission powers, allowing agents to build a comprehensive view of the current state of the network. At time step $t$, each node will broadcast its local state to each of its neighbors and receive their local states to build a local state representation before deciding on the action. The state space is represented by a tuple $\textit{S}=(\mathcal{I},\mathcal{K},\mathcal{Q}^{t},\mathcal{F},\mathcal{L},\{\mathcal{B}_{1},...,\mathcal{B}_{n}\},\{\mathcal{P}_{1},...,\mathcal{P}_{n}\})$, with elements
@@
Action Space. The action space is similarly designed to reflect the decision-making process regarding task offloading, with actions representing the choice of offloading destinations. This setup facilitates the exploration of strategies that balance local processing advantages against the costs and implications of offloading. The action space, $A$, corresponds to the output layer of a DRL agent (see Fig. 2). The action $a_{t}\in\{1,...,N\}$ represents the index of the destination node, which might be one of the neighboring workers or the observed node itself, in case it decides to process the task locally. We use $a_{t}$ interchangeably to denote both the index of the destination node and the act of sending a task to that node, as long as clarity is maintained within the context.
Reward Function. The reward function is a critical component guiding the learning process of the RL agent. It is constructed to reward actions that enhance utility—such as task completions—while penalizing undesirable outcomes like excessive delays or system overloads. This balance encourages the development of nuanced offloading policies that consider various operational constraints and performance metrics. By incorporating reward shaping [27], we further refine the learning process, enabling agents to navigate the complex decision space of task offloading more effectively. This approach not only facilitates faster convergence to optimal policies but also allows for a more nuanced understanding of the trade-offs inherent in Edge Computing task management. The reward shaping term can be defined as $F(s,a,s\textquoteright)=\Xi(s\textquoteright)-\Xi(s)$, where $\Xi(s)$ is a user-defined potential function over a state, $s$ such that the reward becomes $R^{\prime}=R+F$.
@@
where $r_{u}$ is a utility reward and represents the gain over completed tasks, and $\chi_{O}$ is overloading cost weight. Each term of the reward function is explained in detail below. First we introduce an indicator function, $I_{i}(a_{t})$, such that $I_{i}(a_{t})=1$ iff $a_{t}=i$ meaning the task is meant to be processed in node $w_{i}$, otherwise $I_{i}(a_{t})=0$. The reward function for an agent in the context of task offloading is composed of two primary components: the delay function and the cost of overloading. Each incorporates specific parameters and equations that encapsulate the complexities of decision-making in edge-computing environments. The delay function denoted as $D_{i}(s_{t},a_{t})$, is a comprehensive measure that accounts for three distinct temporal aspects associated with task offloading, namely
$D_{i}(s_{t},a_{t})=\chi_{D}^{\text{wait}}T_{i,a_{t}}^{\text{wait}}(\tau_{k})+\chi_{D}^{\text{comm}}T_{i,a_{t}}^{\text{comm}}(\alpha_{k}^{\text{out}})+\chi_{D}^{\text{exc}}T_{i,a_{t}}^{\text{exc}}(\tau_{k}),$
@@
$T_{i,a_{t}}^{\text{wait}}(\tau_{k})=\frac{Q_{i}^{t}}{N^{i}_{\phi}\phi_{i}}+\sum_{j\neq i}\frac{Q_{j}}{N^{j}_{\phi}\phi_{j}}I_{j}(a_{t}),$
where $Q_{i}^{t}$ is the queue size at time $t$, $N^{i}_{\phi}$ and $\phi_{i}$ represent the number of processors and their frequency at node $w_{i}$, respectively. The communication delay $T_{i,a_{t}}^{\text{comm}}(\alpha_{k}^{\text{out}})$, linked to the transmission of $\alpha_{k}^{\text{out}}$ bits of output data, is computed using the Shannon-Hartley theorem. The execution cost difference $T_{i,a_{t}}^{\text{exc}}(\tau_{k})$ between local and target node processing is quantified as
$T_{i,a_{t}}^{\text{exc}}(\tau_{k})=\frac{t\rho_{k}\xi_{k}}{N_{\phi}^{a_{t}}\phi_{a_{t}}}-\frac{t\rho_{k}\xi_{k}}{N_{\phi}^{i}\phi_{i}},$
indicating the variation in processing time due to differences in node capabilities. The cost of overloading, $O_{i}(s_{t},a_{t})$, emphasizes the potential system strain caused by task offloading, expressed as $O_{i}(s_{t},a_{t})=-\log(p_{t}^{Oa_{t}})/3$. The probability of overloading $p_{t}^{O{a_{t}}}$ and the expected queue state $Q^{\prime}_{a_{t}}$ are critical in assessing the impact of offloading decisions on the target node’s workload, calculated as $p_{t}^{O{a_{t}}}=\max\left(0,\frac{Q_{a_{t}}^{\max}-Q_{a_{t}}}{Q_{a_{t}}^{\max}}\right),$ and $Q^{\prime}{a_{t}}=\min(\max(0,Q{a_{t}}-\phi_{a_{t}})+1,Q_{a_{t}}^{\text{max}})$, where $Q_{a_{t}}^{\text{max}}$ is the maximum queue capacity of the offloading target, and $\phi_{a_{t}}$ is its processing rate. These equations collectively frame the decision-making landscape for RL agents, highlighting the intricate balance between task processing efficiency, communication overheads, and system resilience against overloading.
@@ 4.1 System Modeling
Communication Model. Task offloading and reception are simulated under the assumption of a generic wireless communication model, as this is the most common in the literature. Utilizing the Shannon-Hartley theorem [29], we calculate the bits transmitted per second, incorporating factors such as channel bandwidth, channel gain, and transmission power.
In the context of task generation and reception, nodes equipped with a controller protocol play a pivotal role in determining whether to process tasks locally or offload them. This decision-making process is facilitated through the exchange of local state information via one-hop broadcasts. Communication within the network, particularly the offloading interactions among nodes, is predicated on wireless transmission. To quantify the transmission efficiency, specifically the rate at which bits are communicated per second, we invoke the Shannon-Hartley theorem [29]. According to this theorem, the communication latency for transmitting $\alpha$ bits between nodes $w_{i}$ and $w_{j}$ is given by:
$T^{\text{comm}}_{i,j}(\alpha)=\frac{\alpha}{B_{i,j}\log(1+10^{\frac{P_{i}+G_{i,j}-\omega_{0}}{10}})},$ (2)
where $T^{\text{comm}}_{i,j}(\alpha)$ represents the time required for the transmission, $\alpha$ denotes the volume of data in bits, $B_{i,j}$ is the bandwidth of the communication channel between nodes $w_{i}$ and $w_{j}$, $P_{i}$ denotes the transmission power of the source node $w_{i}$, $G_{i,j}$ is the channel gain, and $\omega_{0}$ represents the noise power in the communication channel. This formulation underpins our model for evaluating the communication overhead associated with task offloading in Edge networks.
@@ 4.2 Implementation details
PeersimEnv. PeersimGym integrates the Peersim simulator and the PettingZoo environment (PZ env), facilitating the configuration and execution of simulations through REST requests. This integration is enabled by encapsulating the Peersim simulation within a Spring Boot REST server, providing endpoints for action posting and state retrieval, adhering to PettingZoo standards (Fig 1). A full list of the available configurations can be found in the code repository.
Simulator. At its core, PeersimGym employs the Peersim simulation tool, supporting both event-driven and cycle-driven engines. The simulation models a network as a collection of nodes running various protocols, including client, worker, controller, and a customized Simulation Manager protocol. This Simulation Manager periodically pauses the simulation to process actions and resume, maintaining a cyclical operation that aligns with the specified behaviors of worker, client, and controller protocols.
Through extensive configurability and a focus on customization, PeersimGym presents a powerful tool for simulating Edge network environments and training RL models for task offloading, allowing for more sophisticated and realistic simulations in Edge Computing research.
@@
The main focus of the introduced simulator is on being highly customizable, which includes implementing new protocols that make use of the highly innate modality of the Peersim Tool and the extensions we put in place to manage the simulation of Edge Networks. To tailor the simulation to the required scenario, it is possible to define different protocols for the clients, controllers, and workers, that extend their abstract implementations. Furthermore, different actions and information available to each node can be defined, and the communication model and neighborhood definition can also be customized. We provide multiple base classes that allow doing, out-of-the-box, binary task offloading, where clients generate indivisible tasks and controllers make decisions on where to offload the full tasks; or batch binary task offloading, where clients generate indivisible tasks but controllers decide for each of the tasks arrived in the time after the last offloading decision where they should be processed. More information on what each one does, how to load different modes, and how to create other implementations can be found on the simulator repository. In this paper, we focus on the binary task offloading implementations.

### v3

@@ 1 Introduction
The proliferation of large networks of devices, such as the Internet of Things (IoT), has led to an exponential increase in data generation, requiring significant computational resources for processing. Traditionally, Cloud Computing provided the backbone for such computational demands. However, its limitations in latency and network traffic have become apparent with the growth of device networks [1]. Edge Computing emerged as a paradigm shift, extending the Cloud to bring processing capabilities closer to data sources, mitigating latency and traffic issues. In this evolving landscape, task offloading—the distribution of computational tasks across network participants—has gained prominence, particularly within the Fog and Multi-access Edge Computing (MEC) paradigms. Both paradigms aim to decentralize computing power, bringing it closer to end-users and alleviating the constraints of traditional Cloud Computing [2, 3]. Despite distinctions between MEC and Fog computing, this paper treats them interchangeably, focusing on their shared goal of minimizing device-to-cloud distances [4]. Addressing the challenges of task offloading in such distributed environments involves balancing numerous factors, including task latency, energy consumption, and task completion reliability [5]. Conventional optimization methods often struggle to efficiently manage these complex systems. Multi-Agent Reinforcement Learning (MARL) offers promise in optimizing resource allocation and scheduling to maximize system efficiency or meet specific performance metrics [6, 7]. In the task offloading problem, a reinforcement learning (RL) agent can serve as a decision maker to learn the optimal task-resource allocation strategy by interacting with the environment (i.e., the task and available resources). However, it is impractical to train agents in the real world. Therefore, a standardized, customizable, efficient, and user-friendly simulation tool is essential.
@@
Contributions. Our contributions focus on advancing the field of MARL and Edge computing through the introduction of the PeersimGym environment and comprehensive experimental analysis. The PeersimGym Environment is a highly adaptable simulation platform tailored for the development, training, and evaluation of MARL strategies for task offloading challenges in Edge Computing systems. This environment, built on the synergy between an edge system simulator built with Peersim P2P simulator [8] and the PettingZoo API [9], allows detailed configuration of network topologies, node characteristics, and task parameters, facilitating a wide range of experimental setups. An experimental analysis demonstrates the capability of the PeersimGym to train effective MARL solutions, exemplified by the performance of a Double Deep Q Network and Advantage Actor Critic algorithm. Our analysis compares the performance of DRL agents to several non-RL algorithms across various network configurations and task offloading scenarios, highlighting the advantages of RL approaches in optimizing Edge Computing networks.
@@
The source code for PeersimGym, along with usage documentation and testing resources, is made available in the Simulator repository [^1] https://github.com/FredericoMetelo/peersim-environment, and the Agent repository [^2] https://github.com/FredericoMetelo/TaskOffloadingAgentLibrary, fostering further research and community engagement in advancing Edge Computing solutions.
@@ 3 Enhancing Task Offloading with MARL
To address the complexity of Task Offloading in highly complex edge environments, researchers leverage RL and its subset, DRL, and by exploring the distributed nature of these systems MARL emerges as promising solution. As a tool to develop MARL algorithms, PeersimGym incorporates a Python-based environment tailored for developing, training, and deploying RL models for task offloading, aligned with the PettingZoo framework. With the set of experiments in this paper, we showcase how our simulator can enable further RL and DRL contributions to this field in Sec. 5. Certain Nodes within the network act as RL agents. These agents have the capability to observe the state (albeit partially) and make informed decisions regarding whether to process tasks locally or offload them. This decision-making process is influenced by the need to balance between local processing, the risk of queue overflow, and the costs associated with task offloading, including potential overload of other nodes.
@@
Deep Reinforcement Learning for Task Offloading. Deep Q Networks (DQN) [13] and their variants, such as Double Deep Q Networks (DDQN) [26] and Actor-Critic methods, like the Advantage Actor-Critic (A2C)[12], are prominent DRL approaches applied to task offloading. These methodologies have been shown to stabilize training and enhance learning efficacy through sophisticated neural network architectures that approximate optimal action-selection policies. In PeersimGym, the flexibility of our RL environment supports the integration of various DRL models. For illustrative purposes, we focus on implementations of DDQN and A2C, reflecting their proven effectiveness in recent literature [14, 12]. This choice underscores the potential of DRL to address the complexities associated with task offloading in Edge Computing, as shown in Sec. 5.
Framework and Model Dynamics. Task offloading decisions within our simulated environment are modeled as a Markov Game (MG), accommodating the multi-agent aspect of Edge Computing networks. This formulation extends the Markov Decision Process (MDP) framework to scenarios involving multiple decision-makers, thereby capturing the interactive and competitive nature of task offloading. Single-agent RL is modeled as an MDP, with a sequence of states such that the Markov Property holds, i.e., the next state $s_{t+1}$ depends exclusively on the current state $s_{t}$ and the performed action $a_{t}$. When including multiple agents, most MDP convergence properties do not hold, hence, we formulate our problem as an MG [27]. In our setting, in a network of $N$ nodes, an MG is represented as a tuple $\langle n,\mathcal{S},\mathcal{A},P,R,\gamma\rangle$, where $n$ is the number of agents (nodes with a controller protocol), $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $P$ is an (unknown) transition probability function, $R$ is the reward function, and $\gamma$ is a reward discount factor. Next, we introduce these constructs in more detail.
State Space. The state space in our environment is designed to be highly customizable, enabling the modeling of various Edge Computing scenarios. At time step $t$, each node will broadcast its local state to each of its neighbors and receive their local states to build a local state representation before deciding on the action. The state space is represented by a tuple $\mathcal{S}=(\mathcal{I},\mathcal{K},\mathcal{Q}^{t},\mathcal{F},\mathcal{L},\{\mathcal{B}_{1},...,\mathcal{B}_{n}\},\\ \{\mathcal{P}_{1},...,\mathcal{P}_{n}\})$, with elements
@@
Action Space. The action space is similarly designed to reflect the decision-making process regarding task offloading, with actions representing the choice of offloading destinations. This setup facilitates the exploration of strategies that balance local processing advantages against the costs and implications of offloading. The action space, $\mathcal{A}$, corresponds to the output layer of a DRL agent (see Fig. 2). The action $a_{t}\in\{1,...,N\}$ represents the index of the destination node, which might be one of the neighboring workers or the observed node itself, in case it decides to process the task locally. We use $a_{t}$ interchangeably to denote both the index of the destination node and the act of sending a task to that node, as long as clarity is maintained within the context.
Reward Function. The reward function is a critical component guiding the learning process of the RL agent. It is constructed to reward actions that enhance utility—such as task completions—while penalizing undesirable outcomes like excessive delays or system overloads. This balance encourages the development of nuanced offloading policies that consider various operational constraints and performance metrics. By incorporating reward shaping [28], we further refine the learning process, enabling agents to navigate the complex decision space of task offloading more effectively. This approach not only facilitates faster convergence to optimal policies but also allows for a more nuanced understanding of the trade-offs inherent in Edge Computing task management. The reward shaping term can be defined as $F(s_{t},a_{t},s_{t+1})=\Xi(s_{t+1})-\Xi(s_{t})$, where $\Xi(s_{t})$ is a user-defined potential function over a state, $s_{t}$ such that the reward becomes $R^{\prime}=R+F$.
@@
where $r_{u}$ is a utility reward and represents the gain over completed tasks, and $\chi_{O}$ is overloading cost weight. Each term of the reward function is explained in detail below. First we introduce an indicator function, $I_{i}(a_{t})$, such that $I_{i}(a_{t})=1$ iff $a_{t}=i$, meaning the task is meant to be processed in node $w_{i}$, otherwise $I_{i}(a_{t})=0$. The reward function for an agent in the context of task offloading is composed of two primary components: the delay function and the cost of overloading. Each incorporates specific parameters and equations that encapsulate the complexities of decision-making in edge-computing environments. The delay function denoted as $D_{i}(s_{t},a_{t})$, is a comprehensive measure that accounts for three distinct temporal aspects associated with task offloading, namely
$D_{i}(s_{t},a_{t})=\chi_{D}^{\text{wait}}T_{i,a_{t}}^{\text{wait}}(\tau_{k})+\chi_{D}^{\text{comm}}T_{i,a_{t}}^{\text{comm}}(\alpha_{k}^{\text{out}})+\chi_{D}^{\text{exc}}T_{i,a_{t}}^{\text{exc}}(\tau_{k}),$ (2)
@@
$T_{i,a_{t}}^{\text{wait}}(\tau_{k})=\frac{Q_{i}^{t}}{N^{i}_{\phi}\phi_{i}}+\sum_{j\neq i}\frac{Q_{j}}{N^{j}_{\phi}\phi_{j}}I_{j}(a_{t}),$ (3)
where $Q_{i}^{t}$ is the queue size at time $t$, $N^{i}_{\phi}$ and $\phi_{i}$ represent the number of processors and their frequency at node $w_{i}$, respectively. To quantify the transmission efficiency, specifically the rate at which bits are communicated per second, we invoke the Shannon-Hartley theorem [29]. Accordingly, the communication latency for transmitting $\alpha_{k}^{\text{out}}$ bits between nodes $w_{i}$ and $w_{a_{t}}$ is given by
$T^{\text{comm}}_{i,a_{t}}(\alpha_{k}^{\text{out}})=\frac{\alpha_{k}^{\text{out}}}{B_{i,a_{t}}\log(1+10^{\frac{P_{i}+G_{i,a_{t}}-\omega_{0}}{10}})},$ (4)
where $B_{i,a_{t}}$ is the bandwidth of the communication channel between nodes $w_{i}$ and $w_{a_{t}}$, $P_{i}$ denotes the transmission power of the source node $w_{i}$, $G_{i,a_{t}}$ is the channel gain, and $\omega_{0}$ represents the noise power in the communication channel. This formulation underpins our model for evaluating the communication overhead associated with task offloading in Edge networks. $P_{i}$, $G_{i,a_{t}}$ and $\omega_{0}$ are measured in dB. The execution cost difference $T_{i,a_{t}}^{\text{exc}}(\tau_{k})$ between local and target node processing is quantified as
$T_{i,a_{t}}^{\text{exc}}(\tau_{k})=\frac{\rho_{k}\xi_{k}}{N_{\phi}^{a_{t}}\phi_{a_{t}}}-\frac{\rho_{k}\xi_{k}}{N_{\phi}^{i}\phi_{i}},$ (5)
indicating the variation in processing time due to differences in node capabilities. The cost of overloading, $O_{i}(s_{t},a_{t})$, emphasizes the potential system strain caused by task offloading, expressed as $O_{i}(s_{t},a_{t})=-\log(p_{t}^{{a_{t}}})/3$. The probability of overloading $p_{t}^{{a_{t}}}$ and the expected queue state $Q^{\prime}_{a_{t}}$ are critical in assessing the impact of offloading decisions on the target node’s workload, calculated as $p_{t}^{{a_{t}}}=\max\left(0,\frac{Q_{a_{t}}^{\max}-Q_{a_{t}}}{Q_{a_{t}}^{\max}}\right),$ and $Q^{\prime}_{a_{t}}=\min(\max(0,Q_{a_{t}}-\phi_{a_{t}})+1,Q_{a_{t}}^{\text{max}})$, where $Q_{a_{t}}^{\text{max}}$ is the maximum queue capacity of the offloading target, and $\phi_{a_{t}}$ is its processing rate. These equations collectively frame the decision-making landscape for RL agents, highlighting the intricate balance between task processing efficiency, communication overheads, and system resilience against overloading.
@@ 4.1 System Modeling
Communication Model. Task offloading and reception are simulated under the assumption of a generic wireless communication model, as this is the most common in the literature. Utilizing the Shannon-Hartley theorem [29], we calculate the bits transmitted per second, incorporating factors such as channel bandwidth, channel gain, and transmission power (see Eq. 4).
In the context of task generation and reception, nodes equipped with a controller protocol play a pivotal role in determining whether to process tasks locally or offload them. This decision-making process is facilitated through the exchange of local state information via one-hop broadcasts. Communication within the network, particularly the offloading interactions among nodes, is predicated on wireless transmission.
@@ 4.2 Implementation details
PeersimEnv. PeersimGym integrates the Peersim simulator and the PettingZoo environment (PZ env), facilitating the configuration and execution of simulations through REST requests. This integration is enabled by encapsulating the Peersim simulation within a Spring Boot REST server, providing endpoints for action posting and state retrieval, adhering to PettingZoo standards (Fig. 1). A full list of the available configurations can be found in the code repository.
Simulator. At its core, PeersimGym employs the Peersim simulation tool, supporting both event-driven and cycle-driven engines. The simulation models a network as a collection of nodes running various protocols, including client, worker, controller, and a customized Simulation Manager protocol. This Simulation Manager periodically pauses the simulation to process actions and resume, maintaining a cyclical operation that aligns with the specified behaviors of worker, client, and controller protocols. Through extensive configurability and a focus on customization, PeersimGym presents a powerful tool for simulating Edge network environments and training RL models for task offloading, allowing for more sophisticated and realistic simulations in Edge Computing research.
@@
The main focus of the introduced simulator is on being highly customizable, which includes implementing new protocols that make use of the highly innate modality of the Peersim Tool and the extensions we put in place to manage the simulation of Edge Networks. To tailor the simulation to the required scenario, it is possible to define different protocols for the clients, controllers, and workers, that extend their abstract implementations. Furthermore, different actions and information available to each node can be defined, and the communication model and neighborhood definition can also be customized. We provide multiple base classes that allow doing out-of-the-box, binary task offloading, where clients generate indivisible tasks and controllers make decisions on where to offload the full tasks; or batch binary task offloading, where clients generate indivisible tasks but controllers decide for each of the tasks arrived in the time after the last offloading decision where they should be processed. More information on what each one does, how to load different modes, and how to create other implementations can be found on the simulator repository. In this paper, we focus on the binary task offloading implementations.